OpenStudy (nottim):

Find the absolute maximum and absolute minimum values of the function f(x) = x^4 - 8 x^2 - 10 at interval [-3,-1] and [-4,1].

OpenStudy (nottim):

It says that min is not -17?

OpenStudy (nottim):

but no matter what I do, I keep getting it at -17, which is apparently wrong.

OpenStudy (ikram002p):

f(x) = x^4 - 8 x^2 - 10 f'(x)=4x^3-16x=0 4x(x^2-4)=0 x=0 x=2 x=-2 so its -2 ^^

OpenStudy (nottim):

For the absolute min in both scenarios/intervals?

OpenStudy (nottim):

Oh I see. I made the mistake I always make. Thanks!

OpenStudy (ikram002p):

np